# DPDETR: Decoupled Position Detection Transformer for Infrared-Visible Object Detection

### v1

@@ IV-A Datasets and Evaluation Metric
We conduct experiments on two public infrared-visible object detection datasets: the DroneVehicle Dataset [55] and the KAIST Multispectral Pedestrian Detection Dataset [56]. Both datasets include pairs of RGB-IR images and annotations for performing experiments.
DroneVehicle dataset. The DroneVehicle dataset is a large-scale drone-based infrared-visible oriented object detection dataset. It contains 28,439 pairs of RGB-IR images, covering various scenarios such as urban roads, residential areas, parking lots, and different times from day to night. The dataset provides independently annotated oriented bounding box labels for infrared and visible modalities, covering five categories: car, bus, truck, van, and freight car. To validate the efficacy of our method, we refer to [11] and make some adjustments to the original annotations:
@@
  • Objects for infrared and visible modalities are sorted to ensure they are indexed identically.
The revised dataset consists of 17,990 pairs of images for training and 1,469 pairs for testing.
@@
Learnable Queries PICQS $m$AP
@@
KAIST dataset. The KAIST dataset is a large-scale infrared-visible pedestrian object detection dataset. It contains 95,328 pairs of images captured in driving environments with traffic scenes including streets, campuses, and countryside. Since the original dataset had annotation issues, we use new annotations provided by [14] and inferred using the improved annotations [21]. The training set consists of 7,601 pairs of images, while the test set includes 2,252 frames sampled every 20th frame from videos, among which 1,455 frames are from daytime scenes and 797 frames are from nighttime scenes.
Evaluation metric. For DroneVehicle Dataset, we evaluate detection performance using mean average precision ($m$AP). Specifically, we calculate $m$AP using True Positive (TP) and False Positive (FP) at the Intersection over Union (IoU) threshold of 0.5. For KAIST Dataset, we adopt log average miss rate $MR^{-2}$ over the false positive per image (FPPI) with a range of $[10^{-2},10^{0}]$ to evaluate the pedestrian detection performance. A lower value is better. To further evaluate the performance of the method, we assess its effectiveness under the ‘All’ condition [56] and across six all-day subsets from the test set, which include variations in pedestrian distances and occlusion levels.
@@
Decoder layers num $m$AP Params
@@ IV-B Implementation Details
For the DroneVehicle Dataset, we employ our oriented object detection version, while for the KAIST Dataset, we remove the angle prediction components to achieve horizontal object detection. We select ResNet50 [46] as the backbone of both the infrared and visible branches, with a feature map semantic level of $L$=3. The Decoupled Multispectral Decoder contains six layers and we set the number of attention heads, sampling points, and selected queries as $H$=8, $K$=4, and $N$=300, respectively. During training, we employ basic data augmentations, including random rotation, and flipping, with both training and testing image sizes set to 640 × 640. For the DroneVehicle dataset, we load ResNet50 pre-trained weights and train it for 42 epochs. For the KAIST dataset, we load our base detector pre-trained on the COCO dataset and train it for 24 epochs. The optimizer used is Adam, configured with a learning rate of 0.0001 and a weight decay of 0.0001. All experiments are conducted on an NVIDIA A100 GPU.
@@ IV-C Evaluation on DroneVehicle Dataset
As for infrared and visible oriented object detection, we compare our method with six state-of-the-art single-modality detectors, including Faster R-CNN [1], RetinaNet [48], S2ANet [50], Oriented R-CNN [51], ROI Transformer [52], and Oriented RepPoints [53]. We also make a comparison with 6 infrared-visible multispectral object detection methods, including Halfway Fusion [21], CIAN [54], AR-CNN [14], TSFADet [16], C2Former [12], and GAGTDet [11]. Among them, Halfway Fusion (OBB) is a multispectral oriented object detection method constructed by adding the features of a dual-branch S2ANet. AR-CNN [7], TSFADet [16], and GAGTDet [11] all use both infrared and visible object annotations for constraint. Table I shows the comparison results of these methods. For single-modality detection results, the infrared modality significantly outperforms the visible modality because objects in nighttime scenes are difficult to distinguish from the visible modality. Obviously, infrared-visible multispectral object detection methods are superior to single-modality methods. Our method approaches the SOTA method in the ‘Car’ and ‘Bus’ categories, where the metrics are relatively saturated. It shows significant improvements in the ‘Truck,’ ‘Van,’ and ‘Freight-car’ categories, which are more difficult to distinguish, ultimately surpassing the best method, GAGTDet, by 5.24 $m$AP.
@@
We also provide some detection visualization results on the DroneVehicle Datasets. As shown in Figure 8, C2Former [12] and w/o DP can only output the locations of objects in one modality under misalignment conditions, failing to locate objects in the other modality accurately. The w/o DP even encounters duplicate detection problems due to obvious misalignment situations. In contrast, our DPDETR can accurately locate objects in both modalities and establish the correspondence of the same object in complex misalignment scenarios caused by imaging and motion displacement. DPDETR also achieves better classification and localization by fully leveraging the misaligned object features.
@@ IV-D Ablation Studies
Ablations on Decoupled Position Contrastive DeNoising (DPCDN) training and Query Decoupled Structure (QDS). In Table II, the 1st and 2nd rows display the detection performance of our single modality base detector for infrared and visible modalities, respectively. The 3rd row shows the performance of our base DPDETR on the RGB-IR dual modality. Our base DPDETR outperforms each single modality performance, but it only shows a 1.34 $m$AP improvement over the IR modality due to the challenges in optimizing the decoupled position and correspondence between the two modalities. As shown in the 4th and 5th rows of Table II, we add DPCDN and DQS to achieve an additional improvement of 0.63 $m$AP and 1.52 $m$AP, respectively. Finally, by combining DPCDN and DQS, our method reaches 79.81 $m$AP, representing a 3.75 $m$AP improvement over the IR modality.
@@
Ablations on query initialization. We conduct ablation experiments on query initialization strategies. Specifically, we set three types of decoupled queries as three groups of learnable embeddings for comparison with our Paired IoU-aware Competitive Query Selection (PICQS) initialization method. As shown in Table III, our PICQS method outperforms the learnable queries by 2.33 $m$AP. We attribute this to the complex and varied complementary characteristics of infrared and visible images, thus making our approach of competitively selecting object queries based on image pair characteristics more effective than using pre-learned fixed object queries.
@@ IV-E Effectiveness Analysis of Our Method
Additionally, in Figure 7, we visualize the effectiveness of our Decoupled Position Multispectral Cross-attention in achieving aligned feature fusion under a misalignment scene. In all three decoupled branches, our network correctly focuses on the accurate positions of the same object in both modalities.
@@ IV-F Evaluation on KAIST Dataset
As for infrared and visible pedestrian detection, we remove the angle prediction components of our DPDETR to achieve horizontal object detection. In the experiments, we compare our method with several state-of-the-art multispectral pedestrian detection methods, including ACF [19], Halfway Fusion [21], FusionRPN+BF [20], IAF R-CNN [26], IATDNN+IASS [57], CIAN [54], MSDS-RCNN [28], AR-CNN [14], MBNet [15], TSFADet [16], CMPD [33], GAGTDet [11], and C2Former [12]. As shown in Table IV, our method obtains 26.75 MR, 21.01 MR, and 25.04 MR on the ‘Day,’ ‘Night,’ and ‘All’ conditions under the IoU threshold of 0.5, respectively. Specifically, Our method outperforms the previous best method C2Former [12] by 1.73 MR, 5.66 MR, and 3.35 MR in the ‘Day,’ ‘Night,’ and ‘All’ conditions. Especially, the significant 5.66MR improvement under ‘Night’ conditions indicates that our method is effective in adaptively fusing complementary features and avoiding interference from visible modality features at nighttime.
@@
Methods $m$AP FLOPs Params
TSFADet [16] 73.06 109.8G 104.7M
GAGTDet [11] 74.57 120.6G -
C2Former [12] 74.20 177.7G 120.8M
DPDETR (Ours) 79.81 208.0G 90.1M
We also provide some detection visualization results on the KAIST Dataset. As shown in Figure 9, our DPDETR can accurately locate the pedestrians in both modalities and achieve higher detection confidence even in the case of obvious misalignment scenarios.
IV-G Computational Cost Comparison
Our method is based on Transformers, which typically have a large number of parameters and computational overhead. However, we keep the model’s parameter count and computational cost within a reasonable range by using one-layer Efficient Encoders. As shown in Table VI, compared to other SOTA multispectral object detection methods, our DPDETR does not significantly increase the computational cost (GFLOPs) and even has fewer parameters than these methods.
@@ V Conclusion
In this paper, we propose DPDETR to address the modality misalignment problem by explicitly decoupling and optimizing the object’s position and correspondence in infrared-visible object detection. Through the query decoupled structure, DPDETR effectively focuses on different multispectral complementary features for the object category, visible position, and infrared position, eliminating the optimization gap. In the Position Decoupled Multispectral Deformable Cross-attention module, we achieve accurate adaptive multi-sematic fusion of misaligned features with the constraint of objects’ visible and infrared reference positions. Additionally, our Decoupled Position Contrastive DeNoising Training further enhances DPDETR’s decoupled learning capability. Experiments on drone-vehicle detection and pedestrian detection demonstrate that the proposed method achieves significant improvements compared to other state-of-the-art methods.

### v2

@@ IV-A Datasets and Evaluation Metric
We perform experiments using two publicly available datasets for infrared-visible object detection: the DroneVehicle Dataset [57] and the KAIST Multispectral Pedestrian Detection Dataset [58]. Both datasets include pairs of VIS-IR images and annotations for conducting experiments.
DroneVehicle dataset. The DroneVehicle dataset is a large-scale drone-based infrared-visible oriented object detection dataset. It contains 28,439 pairs of VIS-IR images, covering various scenarios such as urban roads, residential areas, parking lots, and different times from day to night. The dataset provides independently annotated oriented bounding box labels for infrared and visible modalities, covering five categories: car, truck, bus, freight car, and van. To validate the efficacy of our method, we refer to [13] and make some adjustments to the original infrared and visible annotations:
@@
  • Objects in the infrared and visible modalities are organized to ensure consistent indexing.
The adjusted dataset includes 17,990 image pairs for training and 1,469 pairs for testing.
@@
Learnable Queries PICQS mmAP
@@
KAIST dataset. The KAIST dataset is a large-scale infrared-visible pedestrian detection dataset. It contains 95,328 pairs of images captured in driving environments with traffic scenes including streets, campuses, and countryside. Since the original dataset had annotation issues, we use new annotations provided by [16] and inferred using the improved annotations [23]. The training set consists of 7,601 pairs of images, while the test set includes 2,252 pairs, among which 1,455 pairs are from daytime scenes and 797 pairs are from nighttime scenes.
Evaluation metric. For the DroneVehicle Dataset, we evaluate detection performance using mean average precision (mmAP). Specifically, we calculate mmAP using True Positive (TP) and False Positive (FP) at the Intersection over Union (IoU) threshold of 0.5. For the KAIST Dataset, we adopt log average miss rate $MR^{-2}$ over the false positive per image (FPPI) with a range of $10^{-2},10^{0}]$ to evaluate the pedestrian detection performance. A lower value is better. To further evaluate the performance of the method, we assess its effectiveness under the ‘All’ condition [58] and across six all-day subsets from the test set, which include variations in pedestrian distances and occlusion levels.
@@
Decoder layers num mmAP Params
@@ IV-B Implementation Details
For the DroneVehicle Dataset, we employ our oriented object detection version, while for the KAIST Dataset, we remove the angle prediction components to achieve horizontal object detection. We select ResNet50 [48] as the backbone of both the infrared and visible branches, with a feature map semantic level of LL=3. The Decoupled Multispectral Decoder contains six layers and we set the number of attention heads, sampling points, and selected queries as HH=8, KK=4, and NN=300, respectively. During training, we employ basic data augmentations, including random rotation, and flipping, with both training and testing image sizes set to 640 × 640. For the DroneVehicle dataset, we load ResNet50 pre-trained weights and train it for 42 epochs. For the KAIST dataset, we load our base detector pre-trained on the COCO dataset and train it for 24 epochs. The optimizer used is Adam, configured with a learning rate of 0.0001 and a weight decay of 0.0001. All experiments are conducted on an NVIDIA A100 GPU.
@@ IV-C Evaluation on DroneVehicle Dataset
As for infrared and visible oriented object detection, we compare our method with six state-of-the-art single-modality detectors, including Faster R-CNN [1], RetinaNet [50], S2ANet [52], Oriented R-CNN [53], ROI Transformer [54], and Oriented RepPoints [55]. We also make a comparison with 7 infrared-visible multispectral object detection methods, including Halfway Fusion [23], CIAN [56], AR-CNN [16], TSFADet [18], C2Former [14], GAGTDet [13], and DAMSDet [47]. Among them, Halfway Fusion (OBB) is a multispectral oriented object detection method constructed by adding the features of a dual-branch S2ANet. AR-CNN [8], TSFADet [18], and GAGTDet [13] all use both infrared and visible object annotations for constraint. Table I shows the comparison results of these methods. For single-modality detection results, the infrared modality significantly outperforms the visible modality because objects in nighttime scenes are difficult to distinguish from the visible modality. Obviously, infrared-visible multispectral object detection methods are superior to single-modality methods. Our method approaches the SOTA method in the ‘Car’ and ‘Bus’ categories, where the metrics are relatively saturated. It shows significant improvements in the ‘Truck,’ ‘Van,’ and ‘Freight-car’ categories, which are more difficult to distinguish, ultimately surpassing the best method, DAMSDet, by 2.06 mmAP.
@@
We also provide some detection visualization results on the DroneVehicle Datasets. As shown in Figure 8, C2Former [14] and w/o DP can only output the locations of objects in one modality under misalignment conditions, failing to locate objects in the other modality accurately. The w/o DP even encounters duplicate detection problems due to obvious misalignment situations. In contrast, our DPDETR can reliably locate objects in both modalities and establish the correspondence of the same object in complex misalignment scenarios caused by imaging and motion displacement. DPDETR also achieves better classification and localization by fully leveraging the misaligned object features.
@@ IV-D Ablation Studies
Ablations on Decoupled Position Contrastive DeNoising (DPCDN) training and Query Decoupled Structure (QDS). In Table II, the 1st and 2nd rows display the detection performance of our single modality base detector for infrared and visible modalities, respectively. The 3rd row shows the performance of our base DPDETR on the VIS-IR dual modality. Our base DPDETR outperforms each single modality performance, but it only shows a 1.34 mmAP improvement over the IR modality due to the challenges in optimizing the decoupled position and correspondence between the two modalities. As shown in the 4th and 5th rows of Table II, we add DPCDN and DQS to achieve an additional improvement of 0.63 mmAP and 1.52 mmAP, respectively. Finally, by combining DPCDN and DQS, our method reaches 79.81 mmAP, representing a 3.75 mmAP improvement over the IR modality.
@@
Ablations on query initialization. We conduct ablation experiments on query initialization strategies. Specifically, we set three types of decoupled queries as three groups of learnable embeddings for comparison with our Paired IoU-aware Competitive Query Selection (PICQS) initialization method. As shown in Table III, our PICQS method outperforms the learnable queries by 2.33 mmAP. We attribute this to the complex and varied complementary characteristics of infrared and visible images, thus making our approach of competitively selecting object queries based on image pair characteristics more effective than using pre-learned fixed object queries.
@@ IV-E Effectiveness Analysis of Our Method
Additionally, in Figure 7, we visualize the effectiveness of our Decoupled Position Multispectral Cross-attention in achieving aligned feature fusion under a misalignment scene. In all three decoupled branches, our network correctly focuses on the correct positions of the same object in both modalities.
@@ IV-F Evaluation on KAIST Dataset
As for infrared and visible pedestrian detection, we remove the angle prediction components of our DPDETR to achieve horizontal object detection. In the experiments, we compare our method with several state-of-the-art multispectral pedestrian detection methods, including ACF [21], Halfway Fusion [23], FusionRPN+BF [22], IAF R-CNN [28], IATDNN+IASS [59], CIAN [56], MSDS-RCNN [30], AR-CNN [16], MBNet [17], TSFADet [18], CMPD [35], GAGTDet [13], DAMSDet[47], and C2Former [14]. As shown in Table IV, our method obtains 26.75 MR, 21.01 MR, and 25.04 MR on the ‘Day,’ ‘Night,’ and ‘All’ conditions under the IoU threshold of 0.5, respectively. Specifically, Our method outperforms the previous best method C2Former [14] by 1.73 MR, 5.66 MR, and 3.35 MR in the ‘Day,’ ‘Night,’ and ‘All’ conditions. Especially, the significant 5.66MR improvement under ‘Night’ conditions indicates that our method is effective in adaptively fusing complementary features and avoiding interference from visible modality features at nighttime.
@@
Methods mmAP FLOPs Params Speed(fps)
TSFADet [18] 73.06 109.8G 104.7M 18.6
GAGTDet [13] 74.57 120.6G - 17.8
C2Former [14] 74.20 177.7G 120.8M -
DAMSDet (OBB)[47] 78.24 182.7G 77.5M 11.4
DPDETR (Ours) 79.81 208.0G 90.1M 7.2
We also provide some detection visualization results on the KAIST Dataset. As shown in Figure 9, our DPDETR can reliably locate the pedestrians in both modalities and achieve higher detection confidence even in the case of obvious misalignment scenarios.
Figure 10: Visualization of the manually misalignment experiments. The position shift step size along both the x-axis and y-axis is 5 pixels.
IV-G Robustness to the more Serious Misalignment Scenarios
To further evaluate the robustness of the proposed method under more serious modality misalignment problem, we conduct experiments on the DroneVehicle dataset by manually simulating more serious misalignment scenarios. Specifically, we perform both training and evaluation on w/o DP method and DPDETR. In the training phase, we introduce synthetic more serious misalignment by applying position shifts to paired images. Concretely, for each image pair, we randomly select either the infrared or visible image as the Position Shift Image (PSI). The PSI is then shifted along the x-axis by a random offset sampled from the range [13,13][-13,13] pixels. Additionally, with a certain probability, a similar shift is applied along the y-axis to further simulate real-world misalignment conditions. During evaluation, we fix the infrared image and apply a position shift to the visible image by $\Delta x,\Delta y]$ pixels, where $\Delta x,\Delta y\in[-15,15]$. The experimental results are illustrated in Figure 10. It can be seen that our method achieves a mAP of 79.70 under simulated more serious modality misalignment, and the performance degradation of DPDETR on the more serious misaligned validation situations remains minimal. These results demonstrate that our approach exhibits strong robustness to modality misalignment during both training and inference.
IV-H Computational Cost Comparison
As shown in Table VI, compared to other SOTA multispectral object detection methods, our DPDETR does not significantly increase the computational cost (GFLOPs) and even has fewer parameters than these methods. However, since DPDETR simultaneously outputs the locations of objects in both infrared and visible modalities, the inference speed is slightly slower.
@@ V Conclusion
In this paper, we propose DPDETR to address the modality misalignment problem by explicitly decoupling and optimizing the object’s position and correspondence in infrared-visible object detection. Through the query decoupled structure, DPDETR effectively focuses on different multispectral complementary features for the object category, visible position, and infrared position, eliminating the optimization gap. In the Position Decoupled Multispectral Deformable Cross-attention module, we achieve reliable adaptive multi-sematic fusion of misaligned features with the constraint of objects’ visible and infrared reference positions. Additionally, our Decoupled Position Contrastive DeNoising Training further enhances DPDETR’s decoupled learning capability. Experiments on drone-vehicle detection and pedestrian detection demonstrate that the proposed method achieves significant improvements compared to other state-of-the-art methods.